OpenStudy (anonymous):

Find the solution, given the replacement set. 5x - 2y = 6 {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7)

OpenStudy (anonymous):

?

OpenStudy (aroub):

5x-2y=6 {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7) 4(-2)-2(-8)=6 -8+16=6 8=6 5x-2y=6 5(0)-2(-3)=6 0-6=6 -6=6 5x-2y=6 5(1)-2(-1)=6 5+2=6 7=6 5x-2y=6 4(2)-2(4)=6 8-8=6 0=6 5x-2y=6 5(4)-2(7)=6 20-14=6 6=6 So, the answer is (4, 7)